OpenStudy (karabella98):

If a box’s mass is cut in half, but the box is raised to a height four times higher, how does its potential energy change?

OpenStudy (anonymous):

Potential energy: Ep = m * g * h Normal factor: Ep = 1 * g * 1 = g The above conditions show: Ep = 1/2 * g * 4 = 2 g Therefore, the Ep is now twice as large.

OpenStudy (karabella98):

Thank you!